OpenStudy (anonymous):

Given the function f(x) = x + 3x + 4 , find the value of f(–1)

OpenStudy (anonymous):

f(x)=f(-1) f(-1)=-1+3(-1)+4 =-1+2+4 =-1+2=1 1+4=5 F(x)=5

OpenStudy (anonymous):

YOU JUST PLUG IN F(-1) INTO X AND THEN SOLVE HOPE THAT AT LEAST HELPED :)

OpenStudy (anonymous):

f(-1)=-1+3(-1)+4 =-1-3+4 =-4+4 =0

OpenStudy (anonymous):

OOPS MY BAD YUR RIGHT PAT THNX :)

OpenStudy (anonymous):

you are welcome.